# Spacetime Spins: Statistical mechanics for error correction with stabilizer circuits

### v1

@@ IV.3 Logical circuits
Figure 11: Logical error rates for a pair of repetition codes with standard syndrome extraction circuits subjected to a transversal CNOT gate, using both a MWPM decoder and free energy calculations of the spin model, with N=4d2N=4d-2, T=16d1T=16d-1 for MWPM and T=4d1T=4d-1 for the free energy (ML) decoder. The MWPM decoder required a larger circuit depth to observe a threshold; the error rates between these two decoders are therefore not directly comparable. The ML threshold is significantly higher than the MWPM threshold.
@@
We expect the threshold of the model to decrease when subjected to transversal CNOT gates. Figure 11 shows the estimated logical error rates for the standard syndrome extraction circuit, using both MWPM decoding and Monte Carlo simulations of the free energy. A logical error is deemed as occurring if the decoder fails to predict the correct error coset for either of the two code patches. Simulations of the MWPM decoder were unable to produce thresholds for circuit depths smaller than T=16d1T=16d-1, and it was computationally infeasible to simulate spin models for these larger system sizes—for this reason, the logical error rates in Fig. 11 are not directly comparable. For comparison, simulations of the MWPM decoder at T=4d1T=4d-1 (not plotted) consistent with the ML simulations produced a logical error rate around $38\%$ for $d=9$ and $p_{X}=2.9\%$, which is notably larger than the $23\%$ of the ML decoder. Thresholds, on the other hand, are properties of the circuit and noise model, and (up to finite-size effects) remain comparable at different circuit depths. In Fig. 11, we observe an ML threshold at $2.88(2)\%$—lower than the $2.98(1)\%$ of the single code patch. This agrees with the prediction based on the discussion of the defect line introduced by the transversal CNOT. For MWPM, we observe a threshold at $2.68(1)\%$, well below the $2.94(1)\%$ threshold for a single patch. Whereas the difference between ML and MWPM thresholds for the circuits in Sec. IV.2 was around only $0.03$ percentage points, the larger discrepancy seen here may be accounted for by the MWPM algorithm, which requires that each single-qubit error flips at most two detectors (a “graphlike” detector-error model). An $X$ error on a control qubit after the CNOT, however, flips four detectors: two from its own code patch (as usual) and two additional from the target code patch (that now extend across both patches due to the CNOT); an example detector that spans both code patches is shown in Fig. 11. While the MWPM decoder can overcome this by decomposing such errors into products of graphlike errors, doing so introduces suboptimality [higgottSparse2025]. The same considerations apply for the notable gap between the single-patch and CNOT MWPM thresholds.
@@
Instead of one transversal CNOT occurring at the midpoint of the repetition code circuit, we may instead consider the effect of transversal CNOTs occurring within each detector cell of a circuit. Performing numerical simulations (not plotted), the MWPM decoder is now unable to reproduce a threshold, even at extended circuit depths, as an extensive number of non-graphlike errors exist across the detector error model. The ML decoder, on the other hand, sees a threshold near $p_{X}=2.70(5)\%$—a decrease from the $2.88(2)\%$ in Fig. 11. Despite the presence of defects introduced by the transversal CNOTs to every unit cell of the spin model, there nevertheless persists an error-correcting phase at non-zero $p_{X}$.
@@
The discussions in this section naturally extend to other codes with transversal CNOTs, different transversal logic gates, and even lattice surgery or code-switching protocols, for example. Spin models of logical circuits thus allow us to probe the limits of decoding techniques, considering the error-correcting behavior of quantum circuits, rather than merely quantum memories.
@@
Figure 12: Part of a standard syndrome-extraction circuit for the toric code, creating both an $X$-error and $Z$-error detector cell. The code is split into a unit cell with four qubits labeled $Z$, AA, $X$ and $B$, as sketched in Fig. 13. Unit cells are indexed by lattice directions $(x,y,z)$ (with $z$ timelike; separated by dashed lines in the spin diagram). Partially shown CNOTs couple to other unit cells in the directions indicated. The spin diagram for $H_{X}$ shows one unit cell with six spins (white squares and labeled 11-66; spins from adjacent sites are gray) and six bonds [colored to reference Eq. (36)]. Arrows show a bond connecting to another unit cell.
@@ V.1 Circuit compilations
The toric code encodes two logical qubits on a periodic 22D lattice using weight-four $X$ and $Z$ stabilizers on alternating plaquettes [dennisTopological2002, bravyiQuantum1998, kitaevFaulttolerant2003]. Figure 12 shows a standard syndrome-extraction circuit using ancilla qubits [mcewenRelaxing2023, dennisTopological2002]. The corresponding SM model is 33D and has more complex spin geometries than our prior examples. A unit cell for $H_{X}$ is sketched in Fig. 12 and Fig. 13 by applying the elementary spin components from Fig. 3 to the circuit. (The spacetime code is CSS and $H_{Z}$ has an equivalent structure; we focus on $H_{X}$ for simplicity.) Each unit cell at lattice coordinates $(x,y,z)$ has four physical qubits (labeled $Z$, AA, $X$, and $B$), six spins (labeled $\sigma_{i,x,y,z}$ for $i=1,\ldots,6$), and six bonds (signs labeled as $\eta_{j,x,y,z}^{(w)}$ for $j=1,\ldots,6$). [^10] A slightly unusual choice of lattice directions is observed to respect the checkerboard pattern of the plaquettes; to keep $\hat{x}$ and $\hat{y}$ orthogonal we would need to extend to a $12$-spin unit cell. Writing down the interactions from the model, we have a bulk Hamiltonian
@@
with colors matching Fig. 12.111111XX subscripts on $K$ and $\eta$ are omitted for clarity. Compared to the repetition codes, this is a more complex Ising model, with two or four spins per interaction, and weights up to $7$.
Figure 13: A 33D unit cell for the toric code $H_{X}$, based on the syndrome extraction circuit in Fig. 12. Each unit cell is half a detector cell, and has 44 physical qubits, 66 spins (labeled), and 66 bonds (weights unlabeled for simplicity in the diagram).
As we did with the repetition code, we can also consider alternate syndrome-extraction circuits. The wiggling toric code circuit from mcewenRelaxing2023 is shown in Fig. 14. Because ancilla and data qubits are interchanged in the circuit, the second half of each detector cell circuit is the reverse of the first half. The spin model unit cell encompasses both halves and now extends twice as long in time and includes $12$ spins and $12$ bonds. These form the Hamiltonian
@@
Figure 14: Part of a wiggling syndrome-extraction circuit for the toric code, along with its spin diagram for $H_{X}$. Each unit cell again contains four physical qubits $Z$, AA, $X$ and $B$ (cf. Fig. 13), but the timelike direction is twice as long to accommodate the wiggling structure. There are $12$ spins (white squared; spins from adjacent sites are gray) per unit cell and $12$ bonds [colored to reference Eq. (37), and highlighting the symmetric structure of the unit cell].
As with the repetition code, we can form a qualitative understanding of the differences in ML decoding by examining the free energy cost of inserting a spatial logical operator. One representation of the $\bar{X}$ logical is $X$ over all qubits in-between each measurement and reset. On both circuits’ spin diagrams, this flips $\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}\eta^{(3)}_{4,x,y,z}$} and $\color[rgb]{1,0.18,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.18,1}\pgfsys@color@cmyk@stroke{0}{0.82}{0}{0}\pgfsys@color@cmyk@fill{0}{0.82}{0}{0}\eta^{(7)}_{6,x,y,z}$} along all $x$ at some $y,z$. For a ground state of all $\eta=1$ and $\sigma=1$, the energy cost is identical for both circuits. We then consider gauge-equivalent errors. Flipping spin $\sigma_{1,x,y,z}$ deforms the logical error by a gauge operator, [^12] This is the lowest-energy deformation, and therefore the most likely at low temperature (low noise). and changes energy by $\Delta\mathcal{E}_{\text{standard}}=2K^{(5)}$ in the standard circuit and $\Delta\mathcal{E}_{\text{wiggling}}=4K^{(4)}-2K^{(3)}$ in the wiggling circuit. Using Eq. (34), in the low-noise limit we get
@@
These thresholds appear to be considerably lower than the $2.93(2)\%$ threshold for MWPM decoding of a toric code with a phenomenological noise model of bit-flip and measurement errors [wangConfinementHiggs2003]. Our circuit-level noise models are not directly comparable to phenomenological noise models, however, as every detector cell in our circuit supports $44$ spacetime locations where an error can trigger it, compared to six in the phenomenological. A rough comparison can be drawn by considering the threshold as an effective noise rate per detector cell (the probability of a detector being triggered) using Eq. (24): $\frac{1}{2}[1-(1-2\times 0.003655)^{44}]=13.8\%$ for our results, compared to $\frac{1}{2}[1-(1-2\times 0.0293)^{6}]=15.2\%$ for the phenomenological model. These results are more consistent, with a lower threshold potentially due to the increased complexities of error propagation through the circuit’s CNOTs. For further comparison, under circuit-level depolarizing noise the toric code has a MWPM threshold of $0.60\%$ [fowlerHighthreshold2009]. Depolarizing noise often results in higher thresholds due to induced correlations; a toric code with a phenomenological bit-flip noise channel without measurement errors gives a $10.94(2)\%$ ML threshold [honeckerUniversality2001, merzTwodimensional2002], compared to $18.9(3)\%$ for depolarizing noise [bombinStrong2012, ohzekiError2012].
@@ V.2 Gauge symmetries
Gauge symmetries have practical implications for the treatment of the spin models. Gauge-invariant quantities—such as Wilson loops formed by products of spins around closed paths—correspond to observables relevant for determining phases of matter. In the toric code and other topological codes, appropriate Wilson loops may be formed by considering the bonds affected when logical operators are translated in space and time; the complicated bond interactions in our SM model, however, obfuscate the geometric significance of such objects [Eq. (36) and Eq. (37) are not the simple random plaquette gauge theories of phenomenological noise models [dennisTopological2002]].
@@ VI Conclusion
In this work, we established a general framework for mapping stabilizer circuits onto classical statistical mechanics models. Our approach leverages the spacetime code representation of circuits, where the (d+1)(d+1)D circuit becomes a subsystem code in $(d+1)$ spatial dimensions. The probability of any logical error class becomes the partition function of a classical spin Hamiltonian, with gauge operators of the spacetime code corresponding to spin degrees of freedom. This mapping applies to arbitrary noise channels and can be constructed using a spin-diagrams, a graphical language that allows the modular construction of the SM Hamiltonian directly from the circuit elements, and which we described in detail for independent circuit-level $X$-$Z$ noise.
@@
The spin diagram formalism offers both conceptual clarity and practical utility. Each circuit element—idle wires, CNOTs, measurements, resets—contributes a standard building block of spins and interactions. Composing these blocks produces the full SM Hamiltonian without requiring the explicit computation of gauge generators or their commutation relations. These diagrams can be simplified by integrating out low-degree spins, grouping gauge-equivalent errors into weighted interactions while preserving the partition function ratios that determine the phases of the system (error-correcting or non-correcting) and hence maximum-likelihood decoding. The resulting models often take the form of generalized random-bond Ising models, connecting the rich phenomenology of disordered classical systems to the decodability of stabilizer circuits.
@@
Our examples in Sections IV and V illustrate the capabilities of this framework. For the repetition code, we showed that memory and stability experiments map onto spacetime-dual random-bond Ising models on rotated square lattices, providing a clear geometric interpretation of why their thresholds coincide. Different syndrome-extraction circuits, such as a standard ancilla circuit and the wiggling circuit of mcewenRelaxing2023, produce honeycomb lattice models with distinct bond anisotropies. By analyzing the free-energy cost of logical operators in these models, we explained why the standard circuit achieves a slightly higher threshold and lower logical error rates, and confirmed these with MWPM simulations. These results already illustrate how SM models can provide an analytical handle for comparing circuit compilations.
@@
Logical operations also fit naturally within our framework. For example, for the repetition code, we showed how transversal CNOT gates between two code patches introduce defect lines into the SM model, coupling spins across the two lattices and creating higher-order interactions. Monte Carlo simulations revealed that while a single transversal CNOT has minimal impact on the threshold, repeated applications—one per detector cell—reduce the threshold from $2.88(2)\%$ to $2.70(5)\%$. Crucially, an error-correcting phase persists also with these defects—QEC can succeed even when considering evolving logical circuits, and not just static memories—which we not only confirm but also establish its ultimate (i.e., maximum-likelihood decoding) boundaries.
@@
Our framework opens several avenues for future research. An immediate next direction concerns dynamical codes [gottesmanOpportunities2022, derksDynamical2025, sommersDynamically2025, davydovaUniversal2025, setiawanTailoring2025, williamsonDynamical2025, xuFaulttolerant2025, tanggaraStrategic2024, tanggaraSimple2024, fuError2025, derksDesigning2024, kishonyIncreasing2025, shawLowering2025], such as Floquet codes [hastingsDynamically2021, haahBoundaries2022, blackwellCode2025, alamDynamical2024, tangPhases2025, kesselringAnyon2024, claesDynamic2025, rodatzFloquetifying2024, alamDynamical2024, tangPhases2025, sutcliffeDistributed2025, zhangXcube2023, aasenFaultTolerant2023, duaEngineering2024, mclauchlanAccommodating2024, davydovaFloquet2023, motamarriSymTFT2024, moylettLogical2025] or dynamic automorphism codes [davydovaQuantum2024, aitchisonCompeting2025], both in platforms with native two-qubit measurements [grans-samuelssonImproved2024, grans-samuelssonFaulttolerant2025], and in other systems where one can again compare distinct circuit compilations.
@@
For these and further problems, one can leverage the systematic and modular nature of spin diagrams which makes them amenable to automation: given a circuit description, one could algorithmically generate the corresponding Hamiltonian and estimate thresholds via Monte Carlo methods. This would complement existing approaches for comparisons of circuit compilations, gate orderings, or scheduling strategies for a given code. It may be fruitful to also consider other circuits, such as different compilations of the surface code syndrome-extraction circuit, surface code transversal logic gates or lattice surgery, magic state factories, or code-switching architectures. There has been recent progress in developing syndrome-extraction circuits that tolerate qubit dropouts [debroyLUCI2024, wolanskiAutomated2025, strikisQuantum2023, ankerOptimized2025]; it would be interesting to see how these modified circuits affect the spin models—if lost qubits manifest as defects, and if so, what are their properties and effect on the phases of the SM models? Through our framework, these dynamical perspective can now all be readily understood through the perspective of statistical mechanics.
@@
Our framework also opens conceptual directions for exploration. One of these concerns mixed-state phases of matter [lyonsUnderstanding2024, baoMixedstate2023, liReplica2024, fanDiagnostics2024, sohalNoisy2025, sangMixedState2024, rakovszkyDefining2024, luoTopological2025, chenSeparability2024, lessaStrongtoWeak2025, leeMixedState2025], providing an alternative viewpoint on noise-resilient phases. This approach can also relate the decodability of a quantum code to classical SM systems, but instead of spins enumerating equivalent error strings consistent with a syndrome, they encode the syndrome itself when a codestate is subjected to an error channel. These mixed-state models are Kramers-Wannier duals to the SM models discussed in this paper, and the mapping from spacetime codes discussed here would follow analogously. The resulting models will describe a static mixed state in $(d+1)$ spatial dimensions—a counterpart of the time-evolving error-corrupted physical system in $d$ spatial dimensions. (See also Ref. [negariSpacetime2024] for a mixed-state approach to fault tolerance via static (d+1)(d+1)D states.) It would be interesting to study how the corresponding information-theoretic quantities can capture features accounted for by our SM models.

### v2

@@ IV.3 Logical circuits
Figure 12: Logical error rates for a pair of repetition codes with standard syndrome extraction circuits subjected to one transversal CNOT gate, using a MWPM decoder and free energy calculations, with N=4d2N=4d-2, T=16d1T=16d-1 for MWPM and T=4d1T=4d-1 for the free energy (ML) decoder. (The MWPM decoder required a larger circuit depth to observe a threshold, cf. Fig. A.1; the error rates between these two decoders are therefore not directly comparable.) The ML threshold is significantly higher than the MWPM threshold.
Figure 13: Logical error rates for a pair of repetition codes with standard syndrome extraction circuits subjected to $d$ transversal CNOT gates, using both a MWPM decoder and free energy calculations, with N=4d2N=4d-2, T=4d1T=4d-1 for both. A threshold persists for only the free energy (ML) decoder.
@@
We expect the threshold of the model to decrease when subjected to transversal CNOT gates. Figure 12 shows the estimated logical error rates for the standard syndrome extraction circuit, using both MWPM decoding and Monte Carlo simulations of the free energy. A logical error occurs if the decoder fails to predict the correct error coset for either of the two code patches. Simulations of the MWPM decoder were unable to produce thresholds for circuit depths smaller than T=16d1T=16d-1, and it was computationally infeasible to simulate spin models for these larger system sizes—for this reason, the logical error rates in Fig. 12 are not directly comparable. Simulations of the MWPM decoder at T=4d1T=4d-1 (cf. Fig. A.1) consistent with the ML simulations produced a logical error rate around $36\%$ for $d=9$ and $p_{X}=2.8\%$, which is notably larger than the $23\%$ of the ML decoder. Thresholds, on the other hand, are properties of the circuit and noise model, and (up to finite-size effects) remain comparable at different circuit depths. In Fig. 12, for one transversal CNOT, we observe an ML threshold at $2.92(2)\%$—lower than the $3.03(1)\%$ without the gate. For MWPM, we observe a threshold at $2.68(1)\%$, well below the $2.94(1)\%$ threshold without the gate.
@@
Instead of one transversal CNOT occurring at the midpoint of the repetition code circuit, we may also consider CNOTs occurring within each detector cell of a circuit. Performing numerical simulations (cf. Fig. 13), the MWPM decoder is now unable to produce a threshold, even at extended circuit depths (cf. Fig. A.1). The ML decoder, on the other hand, has a threshold at $p_{X}=2.69(2)\%$. Despite the presence of defects introduced by the transversal CNOTs to every unit cell of the spin model, there nevertheless persists an error-correcting phase at non-zero $p_{X}$.
Whereas the difference between ML and MWPM thresholds for the circuits in Sec. IV.2 was around only $0.09$ percentage points, the larger discrepancy seen here may be accounted for by the MWPM algorithm, which requires that each single-qubit error flips at most two detectors (a “graphlike” detector-error model). An $X$ error on a control qubit after the CNOT, however, flips four detectors: two from its own code patch (as usual) and two additional from the target code patch (that now extend across both patches due to the CNOT); an example detector that spans both code patches is shown in Fig. 11. While the MWPM decoder can overcome this by decomposing such errors into products of graphlike errors, doing so introduces suboptimality [63]. The same considerations apply for the notable gap between the with-CNOT and without-CNOT MWPM thresholds.
@@
The discussions in this section naturally extend to other codes with transversal CNOTs, different transversal logic gates, lattice surgery, or code-switching protocols. Spin models of logical circuits thus allow us to probe the limits of decoding techniques, considering the error-correcting behavior of quantum circuits, rather than merely quantum memories.
@@
Figure 14: Part of a standard syndrome-extraction circuit for the toric code, creating both an $X$-error and $Z$-error detector cell. The code’s unit cell has four physical qubits $Z$, AA, $X$ and $B$, cf. Fig. 15, and is indexed by lattice directions $(x,y,z)$ (with $z$ timelike; separated by dashed lines in the spin diagram). Partially shown CNOTs couple to other unit cells in the directions indicated. The spin diagram for $H_{X}$ shows one unit cell with six spins (white squares and labeled 11-66; spins from adjacent sites are gray) and six bonds [colored to reference Eq. (36)]. Arrows connect bonds to other unit cells.
@@ V.1 Circuit compilations
The toric code encodes two logical qubits on a periodic 22D lattice using weight-four $X$ and $Z$ stabilizers on alternating plaquettes [37, 21, 71]. Figure 14 shows a standard syndrome-extraction circuit using ancilla qubits [84, 37]. The corresponding SM model is 33D and has more complex spin geometries than our prior examples. A unit cell for $H_{X}$ is sketched in Fig. 14 and Fig. 15 by applying the elementary spin components from Fig. 4 to the circuit. (The spacetime code is CSS and $H_{Z}$ has an equivalent structure; we focus on $H_{X}$ for simplicity.) Each unit cell at lattice coordinates $(x,y,z)$ has four physical qubits (labeled $Z$, AA, $X$, and $B$), six spins (labeled $\sigma_{i,x,y,z}$ for $i=1,\ldots,6$), and six bonds (signs labeled as $\eta_{j,x,y,z}^{(w)}$ for $j=1,\ldots,6$). Writing down the interactions from the model, we have a bulk Hamiltonian
@@
with colors matching Fig. 14 ($X$ subscripts on $K$ and $\eta$ are omitted for clarity). Compared to the repetition codes, this is a more complex Ising model, with two or four spins per interaction, and weights up to $7$.
Figure 15: A 33D unit cell for the toric code $H_{X}$, based on the syndrome extraction circuit in Fig. 14. Each unit cell is half a detector cell, and has 44 physical qubits, 66 spins (labeled), and 66 bonds (weights unlabeled for simplicity in the diagram). The lattice directions follow the checkerboard pattern of the plaquettes; to keep $\hat{x}$ and $\hat{y}$ orthogonal we would need to extend to a $12$-spin unit cell.
As we did with the repetition code, we can also consider alternate syndrome-extraction circuits. The wiggling toric code circuit from McEwen et al. [84] is shown in Fig. 16. Because ancilla and data qubits are interchanged in the circuit, the second half of each detector cell circuit is the reverse of the first half. The spin model unit cell encompasses both halves and now extends twice as long in time and includes $12$ spins and $12$ bonds. These form the Hamiltonian
@@
Figure 16: Part of a wiggling syndrome-extraction circuit for the toric code, along with its spin diagram for $H_{X}$. Each unit cell again contains four physical qubits $Z$, AA, $X$ and $B$ (cf. Fig. 15), but the timelike direction is twice as long to accommodate the wiggling structure. There are $12$ spins (white squares; spins from adjacent sites are gray) per unit cell and $12$ bonds [colored to reference Eq. (37), and highlighting the symmetric structure of the unit cell].
As with the repetition code, we can form a qualitative understanding of the differences in ML decoding by examining the free energy cost of inserting a spatial logical operator. One representation of the $\bar{X}$ logical is $X$ on qubits along a non-contractible cycle of the lattice, in-between each measurement and reset. On both circuits’ spin diagrams, this flips $\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}\eta^{(3)}_{4,x,y,z}$} and $\color[rgb]{1,0.18,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.18,1}\pgfsys@color@cmyk@stroke{0}{0.82}{0}{0}\pgfsys@color@cmyk@fill{0}{0.82}{0}{0}\eta^{(7)}_{6,x,y,z}$} along all $x$ at some $y,z$. For a ground state of all $\eta=1$ and $\sigma=1$, the energy cost is identical for both circuits. We then consider gauge-equivalent errors. Flipping spin $\sigma_{1,x,y,z}$ deforms the logical error by a gauge operator, [^10] This is the lowest-energy deformation, and therefore the most likely at low temperature (low noise). and changes energy by $\Delta\mathcal{E}_{\text{standard}}=2K^{(5)}$ in the standard circuit and $\Delta\mathcal{E}_{\text{wiggling}}=4K^{(4)}-2K^{(3)}$ in the wiggling circuit. Using Eq. (34), in the low-noise limit we get
@@
These thresholds appear to be considerably lower than the $2.93(2)\%$ threshold for MWPM decoding of a toric code with a phenomenological noise model of bit-flip and measurement errors [119]. Our circuit-level noise models are not directly comparable to phenomenological noise models, however, as every detector cell in our circuit supports $44$ spacetime locations where an error can trigger it, compared to six in the phenomenological. A rough comparison can be drawn by considering the threshold as an effective noise rate per detector cell (the probability of a detector being triggered) using Eq. (24): $\frac{1}{2}[1-(1-2\times 0.003655)^{44}]=13.8\%$ for our results, compared to $\frac{1}{2}[1-(1-2\times 0.0293)^{6}]=15.2\%$ for the phenomenological model. These results are more consistent, with a lower threshold potentially due to the increased complexities of error propagation through the circuit’s CNOTs. For further comparison, under circuit-level depolarizing noise the toric code has a MWPM threshold of $0.60\%$ [48]. Depolarizing noise often results in higher thresholds due to induced correlations; a toric code with a phenomenological bit-flip noise channel without measurement errors gives a $10.94(2)\%$ ML threshold [64, 86], compared to $18.9(3)\%$ for depolarizing noise [17, 90].
@@ V.2 Gauge symmetries
Gauge symmetries have practical implications for the treatment of the spin models. Gauge-invariant quantities—such as Wilson loops formed by products of spins around closed paths—correspond to observables relevant for determining phases of matter. In the toric code and other topological codes, appropriate Wilson loops may be formed by considering the bonds affected when logical operators are translated in space and time; the complicated bond interactions in our SM model, however, obfuscate the geometric significance of such objects [Eq. (36) and Eq. (37) are not the simple random plaquette gauge theories of phenomenological noise models [37]].
@@ VI Conclusion
In this work, we established a general framework for mapping stabilizer circuits onto classical statistical mechanics models. Our approach leverages the spacetime code representation of circuits, where the (d+1)(d+1)D circuit becomes a subsystem code in $(d+1)$ spatial dimensions. The probability of any logical error class becomes the partition function of a classical spin Hamiltonian, with gauge operators of the spacetime code corresponding to spin degrees of freedom. This mapping applies to arbitrary noise channels and can be visualized using spin diagrams, which we described in detail for independent circuit-level $X$-$Z$ noise.
@@
The spin diagram formalism is a modular graphical language that encodes gauge-equivalent errors and their relative probabilities, and offers both conceptual clarity and practical utility. Each circuit element—idle wires, CNOTs, measurements, resets—contributes a standard building block of spins and interactions that combine to create the SM Hamiltonian without requiring explicit computation of gauge generators or their commutation relations. These diagrams can be simplified by integrating out low-degree spins, grouping gauge-equivalent errors into weighted interactions while preserving the partition function ratios that determine the phases of the system (error-correcting or non-correcting) and hence maximum-likelihood decoding. The resulting models often take the form of generalized random-bond Ising models, connecting the rich phenomenology of disordered classical systems to the decodability of stabilizer circuits.
@@
Our examples in Sections IV and V illustrate the capabilities of this framework. For the repetition code, we showed that memory and stability experiments map onto spacetime-dual random-bond Ising models on rotated square lattices, providing a clear geometric interpretation of why their thresholds coincide. Different syndrome-extraction circuits, such as a standard ancilla circuit and the wiggling circuit of McEwen et al. [84], produce honeycomb lattice models with distinct bond anisotropies. By analyzing the free-energy cost of logical operators in these models, we explained why the standard circuit achieves a slightly higher threshold and lower logical error rates, and confirmed these with MWPM and Monte Carlo free energy simulations. These results already illustrate how SM models can provide an analytical handle for comparing circuit compilations.
@@
Logical operations also fit naturally within our framework. For example, for the repetition code, we showed how transversal CNOT gates between two code patches introduce defect lines into the SM model, coupling spins across the two lattices and creating higher-order interactions. Monte Carlo simulations revealed that while a single transversal CNOT has minimal impact on the threshold, repeated applications—one per detector cell—reduce the threshold from $2.92(2)\%$ to $2.69(2)\%$. Crucially, an error-correcting phase persists also with these defects—QEC can succeed even when considering evolving logical circuits, and not just static memories—which we not only confirm but also establish its ultimate (i.e., maximum-likelihood decoding) boundaries.
@@
Our framework opens several avenues for future research. An immediate next direction concerns dynamical codes [57, 39, 106, 32, 102, 122, 125, 113, 112, 49, 38, 69, 103], such as Floquet codes [61, 60, 16, 5, 111, 68, 31, 96, 5, 111, 109, 128, 1, 40, 85, 34, 87, 88] or dynamic automorphism codes [33, 4], both in platforms with native two-qubit measurements [59, 58], and in other systems where one can again compare distinct circuit compilations. It may be fruitful to also consider other circuits, such as different compilations of the surface code syndrome-extraction circuit, surface code transversal logic gates or lattice surgery, magic state factories, or code-switching architectures. There has been recent progress in developing syndrome-extraction circuits that tolerate qubit dropouts [35, 123, 108, 6]; it would be interesting to see how these modified circuits affect the spin models—if lost qubits manifest as defects, and if so, what are their properties and effect on the phases of the SM models? Through our framework, these dynamical perspectives can now all be readily understood through the perspective of statistical mechanics.
For these and further problems, one can leverage the systematic and modular nature of spin diagrams, which makes them amenable to automation: given a circuit description, one could algorithmically generate the corresponding Hamiltonian and estimate thresholds via Monte Carlo methods. This would complement existing approaches for comparisons of circuit compilations, gate orderings, or scheduling strategies for a given code. Spin diagrams fundamentally encode gauge-equivalent error configurations, and by considering dominant contributions to the partition function, it may also be possible to form analytic bounds on the entropic contributions to logical failure rates of circuits [14, 15].
@@
Our framework also opens conceptual directions for exploration. One of these concerns mixed-state phases of matter [81, 9, 78, 47, 105, 100, 95, 80, 26, 76, 75], providing an alternative viewpoint on noise-resilient phases. This approach can also relate the decodability of a quantum code to classical SM systems, but instead of spins enumerating equivalent errors consistent with a syndrome, they encode the syndrome itself when a codestate is subjected to a noise channel. These mixed-state models are Kramers-Wannier duals to the SM models discussed in this paper, and the mapping from spacetime codes discussed here would follow analogously. The resulting models will describe a static mixed state in $(d+1)$ spatial dimensions—a counterpart of the time-evolving error-corrupted physical system in $d$ spatial dimensions. (See also Ref. [89] for a mixed-state approach to fault tolerance via static (d+1)(d+1)D states.) It would be interesting to study how the corresponding information-theoretic quantities can capture features accounted for by our SM models.